OpenStudy (cheesecakekitten):

Which ordered pair is on the graph of the equation 3x + 7y = 11? A. (1, -2) B. (6, -1) C. (0, -4) D. (0, 4)

OpenStudy (alexandervonhumboldt2):

substitute each point into the equation and check which will be correct

OpenStudy (cheesecakekitten):

POOJA ONLY

pooja195 (pooja195):

Plug in the given coordinates and see which makes atrue statement like this: 3(3)+7(-2)=11 Is this true? or false?

OpenStudy (alexandervonhumboldt2):

why? i haven't done anything bad.

OpenStudy (shiburinga):

She just prefers one person, probably less confusing. Lets go help other people, Alex :D

pooja195 (pooja195):

\[\huge~\rm~3(3)+7(-2)=11\]

OpenStudy (cheesecakekitten):

Just a sec

OpenStudy (cheesecakekitten):

False

pooja195 (pooja195):

It was supposed to be 1,-2 >.< sorry! \[\huge~\rm~3(1)+7(-2)=11\] But tis still false so eh doesnt matter :P Ok next one 3(6)+7(−1)=11 True or false?

OpenStudy (cheesecakekitten):

True

pooja195 (pooja195):

Good so we know that, that is the answer :)

OpenStudy (cheesecakekitten):

Thanks a thousand Pooj.

pooja195 (pooja195):

No probs :)